# HalluciBot: Is There No Such Thing as a Bad Question?

### v1

@@ 7 Conclusion
We propose a heretofore unexplored realm of hallucination mitigation - predicting a query’s hallucination probability. HalluciBot empirically estimates how the query itself may induce hallucination. HalluciBot’s training corpus consists of diverse scenarios and domains to ensure robustness; we score an accuracy level of 73% for binary and top three multi-class labels. In essence, HalluciBot lays the foundation for a reproducible method of hallucination prediction that absorbs the cost of iterative generations during training. Moreover, the training process of HalluciBot delivers insight into the performance of gpt-3.5-turbo on lexical perturbations for 13 datasets - and present discoveries on how perturbations may induce hallucinations. Practically, institutions can implement HalluciBot to measure user accountability - penalizing those who input queries that have a hallucination probability above a certain threshold. Thus, HalluciBot’s academic and practical contributions add to the ever-growing concerted effort of enabling a robust language generation ecosystem for society.
Limitations
Despite the promising results of HalluciBot, there are limitations to be acknowledged. Firstly, our approach relies on automated LLM crowdsourcing to generate query variations, which can introduce a certain degree of variability and noise in the data. Although we took measures to ensure the quality of LLM-crowdsourced data through examining well-formedness, there may be bias or error amongst variations. Secondly, our approach is computationally expensive during training, requiring multiple LLM agents and generating $n$ variations per query. Thirdly, our datasets are English-centric, so our experimental results for robustness and consistency can only be interpreted and applied to English. Thus, it may be difficult to conclude if HalluciBot could withstand stress-testing if every query was perturbed into a new language, given that our BERT and RoBERTa backbones are English-centric. Fourthly, HalluciBot is trained soley on the outputs of gpt-3.5-turbo; however, as the framework is adaptable, future studies can train HalluciBot on any mixture of LLMs. Finally, there may be inherent bias in using gpt-3.5-turbo as both the Query Perturbator and the Output Generator. However, even though there is significant agreement regardless of correctness (Table 17), we feel confident that this is not a limiting factor of our work. Since every label is observed in significant quantities, we can assume that perturbations helped decouple semantic bias from gpt-3.5-turbo’s general knowledge.
@@
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ A.1 What is Extractive QA?
Extractive Question-Answering involves extracting answers directly from a given context. It can be accomplished using span-detection encoders that predict the start and end tokens of the relevant portion in the context (Devlin et al., 2018). Another approach is Retrieval-Augmented Generation (RAG), where the model generates the answer based on a selected passage containing the necessary information (Lewis et al., 2021; Guu et al., 2020).
@@ A.2 What is Multiple Choice QA?
Multiple Choice Question-Answering is a task in few-shot learning where a model is given a question and a set of answer choices. Usually, one of the choices is correct, while the rest are distractors. In an encoder-based approach, the model replicates the question for each choice, generates a scalar response, and applies a Softmax across the choices to determine the best answer (Devlin et al., 2018). In a generative setting, the model can generate an answer choice instead of selecting from the given options, using few-shot prompting techniques (Wei et al., 2022).
@@ A.3 What is Abstractive QA?
Abstractive Question-Answering refers to the generation of answers without access to context or candidates. Also known as Closed-book Generative QA, where no context is provided, abstractive methods are used to generate answers based solely on the question (Wei et al., 2022). Encoder models can be employed in this approach, which convert commonly occurring answer tokens into class labels (Kim et al., 2021). A multi-class model is then trained using Softmax regression to predict the answer from the question.
@@ A.4 What is Temperature and Nucleus Sampling?
Greedy decoding takes the most likely predicted token of the output distribution per sequence step, but often yields poor results and a lack of diversity. To combat this, alternative approaches leverage multinomial sampling across the distribution to sample less likely tokens. Temperature $T$ is used to control the smoothing of the distribution. As $T\to\infty$, the output distribution is smoother and more uniform; therefore unlikely tokens become more likely to sample. As $T\to 0$, the distribution approaches a Kronecker Delta ($\delta_{ij}$) function, centered on the token with the most probability mass (and mimics greedy decoding strategies). Temperature is often implemented as an adjusted Softmax function (Holtzman et al., 2020).
@@
However, sampling across the full vocabulary can lead an LLM to produce extremely unlikely generations. To restrict the space of possible outcomes, two approaches exist: Top-$K$ Sampling and Nucleus Sampling. Top-$K$ sampling is implemented to include tokens that are in the Top-$K$ most probable tokens (Fan et al., 2018; Holtzman et al., 2018; Radford et al., 2018). One limitation is the static window of possible tokens considered, where unlikely tokens can still be considered even if most of the probability mass is distributed amongst fewer tokens than $K$.
Nucleus Sampling, instead of considering the Top-$K$ most probable tokens in vocabulary $V$, only considers the smallest set of tokens that have a cumulative likelihood greater than some threshold $p_{i}$. Therefore, each decoding step will consider the most likely Top-$p_{i}$ tokens $V^{(p_{i})}\subset V$, automatically eliminating improbable tokens from being accidentally sampled through masking (Holtzman et al., 2020).
@@ A.5 What is a Multi-Agent Simulation?
In a Multi-Agent Simulation, independent agents collaborate and interact to find a solution (Russell and Norvig, 2009). In the context of our experiments, we have two types of agents: the Query Perturbator and $n+1$ Output Generators. The purpose of the Multi-Agent Simulation is to generate independent outputs after perturbing the query $q_{0}$ to observe the rate of hallucination. The presence of $n+1$ queries and agents is to ensure a balanced representation of hallucination, preventing skewed results towards either extreme. The set of outputs generated by the agents is then evaluated against the ground truth value $y$ to estimate the hallucination rate. Through multiple samplings across all 369,837 queries, we train HalluciBot to accurately identify and assess the risk of hallucination.
@@ A.6 What is Monte Carlo Sampling?
Monte Carlo sampling is a technique used to approximate unknown quantities of interest when it’s difficult to derive an exact solution (Murphy, 2012). This is particularly useful when dealing with hidden latent variables that cannot be directly observed, such as the case of hallucination caused by interacting with a complex random variable (e.g., a Large Language Model). To address this challenge, we use Monte Carlo sampling. The idea is to conduct a Multi-Agent Simulation for each query in our training set. By doing so, we can estimate the probability of hallucination by observing the accuracy rate per query across lexical perturbations. The more simulations we conduct, the more accurate our estimation becomes. In our experiments, we sample from a series of 369,837 Multi-Agent Simulations that examine hallucination rates across perturbations for any query $q_{0}$. Therefore, we can estimate the hallucination rate $p_{h}(q_{0})$ for query $q_{0}$, given the ground truth $y$ and set of answers $a_{i}\in\mathcal{A}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\big{[}a_{i}\not=y\big{]}$
@@ A.7 What is Ordinal Regression?
Expanding to several classes $K$, we can divide the probability space into $K-1$ cutpoints $c_{1},\dots,c_{K-1}$ with the property that $c_{1}<c_{2}<\cdots c_{K-1}$, such that the probability for each class is defined as follows (Winship and Mare, 1984):
@@
To enforce the property $c_{1}<c_{2}<\cdots c_{K-1}$ while allowing our thresholds to be differentiable, we employ a cumulative sum on a set of unbounded, learnable parameters $\theta_{1},\dots,\theta_{K-1}$, transformed by a Softplus ($\beta=1$) to avoid adding negative values; this ensures our thresholds are always increasing (Zheng et al., 2015). Therefore, each cutpoint $c_{k}$ will be of the from:
@@
Appendix B Additional Experiment Details
@@ B.1 LLM Agent Settings
B.2 Token Usage & Statistics
@@
B.3 HalluciBot Training Setup & Metrics
The parameters to our AWS EC2 training instance are provided in Table 7. Additionally, we report our training metrics, including GPU hours and total floating point operations (FLOS), in Table 9. Finally, the parameters for our backbones are enumerated in Table 10.
@@ B.3 HalluciBot Training Setup & Metrics
Model Size (Hours) FLOS Steps / Second / Second Loss
@@
Model Size (Hours) FLOS Steps / Second / Second Loss
@@
Table 9: HalluciBot training statistics for our binary and multi-class experiments. Size is the number of learnable parameters in the model. Total FLOS is the total number of floating-point operations conducted in the model. Update Steps is the number of parameter updates the Adam optimizer (Kingma and Ba, 2017) performs on the model. We trained each model for 5 epochs in total. Training parameters can be found in Table 10. There is no difference in FLOS between BERT and RoBERTa base models as RoBERTa’s parameter increase is concentrated in a larger vocabulary.
@@
Table 10: Model configuration and hyper-parameters for our HalluciBots. Configurations changed from defaults highlighted in bold. Note that for training batch size and gradient accumulation steps, we only altered roberta-large’s hyper-parameters to fit onto our GPU. All models were trained using HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba, 2017).

### v3

@@ 6 Conclusion
We propose a heretofore relatively unexplored realm of hallucination mitigation - predicting a query’s hallucination probability. HalluciBot empirically estimates how the query itself may induce hallucination and its training corpus consists of diverse scenarios and domains to ensure robustness. Institutions can implement HalluciBot to measure user accountability and improve the robustness of LLM’s performance via our H4Rs (“Ratiocinate”, “Rewrite”, “Rank”, “Route”). Thus, HalluciBot’s academic and practical contributions add to the ever-growing concerted effort of enabling a robust language generation ecosystem for society.
Limitations. HalluciBot relies on automated LLM crowdsourcing via Monte Carlo sampling to generate perturbations, which may introduce noise. Sampling is computationally expensive during training, but is balanced by its ability to curate large, diverse corpora spanning various domains and scenarios. HalluciBot can be trained on a mixture of LLMs or used as a proxy-reward model for RL-tuned generators.
@@
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ A.1 What is Extractive QA?
Extractive Question-Answering involves extracting answers directly from a given context. It can be accomplished using span-detection encoders that predict the start and end tokens of the relevant portion in the context (Devlin et al. 2018). Another approach is Retrieval-Augmented Generation (RAG), where the model generates the answer based on a selected passage containing the necessary information (Lewis et al. 2020; Guu et al. 2020).
@@ A.2 What is Multiple Choice QA?
Multiple Choice Question-Answering is a task in few-shot learning where a model is given a question and a set of answer choices. Usually, one of the choices is correct, while the rest are distractors. In an encoder-based approach, the model replicates the question for each choice, generates a scalar response, and applies a Softmax across the choices to determine the best answer (Devlin et al. 2018). In a generative setting, the model can generate an answer choice instead of selecting from the given options, using few-shot prompting techniques (Wei et al. 2022).
Figure 7: To assess the possibility of hallucination, existing literature focuses on either the outputs or the intermediate outputs within the model. Our method instead focuses on directly assessing the quality of the query, which is quantified as as how likely it will lead to a hallucination.
@@ A.3 What is Abstractive QA?
Abstractive Question-Answering refers to the generation of answers without access to context or candidates. Also known as Closed-book Generative QA, where no context is provided, abstractive methods are used to generate answers based solely on the question (Wei et al. 2022). Encoder models can be employed in this approach, which convert commonly occurring answer tokens into class labels (Kim, Son, and Kim 2021). A multi-class model is then trained using Softmax regression to predict the answer from the question.
@@ A.4 What is Temperature and Nucleus Sampling?
Greedy decoding takes the most likely predicted token of the output distribution per sequence step, but often yields poor results and a lack of diversity. To combat this, alternative approaches leverage multinomial sampling across the distribution to sample less likely tokens. Temperature $T$ is used to control the smoothing of the distribution. As $T\to\infty$, the output distribution is smoother and more uniform; therefore unlikely tokens become more likely to sample. As $T\to 0$, the distribution approaches a Kronecker Delta ($\delta_{ij}$) function, centered on the token with the most probability mass (and mimics greedy decoding strategies). Temperature is often implemented as an adjusted Softmax function (Holtzman et al. 2020).
@@
However, sampling across the full vocabulary can lead an LLM to produce extremely unlikely generations. To restrict the space of possible outcomes, two approaches exist: Top-$K$ Sampling and Nucleus Sampling. Top-$K$ sampling is implemented to include tokens that are in the Top-$K$ most probable tokens (Fan, Lewis, and Dauphin 2018; Holtzman et al. 2018; Radford et al. 2018). One limitation is the static window of possible tokens considered, where unlikely tokens can still be considered even if most of the probability mass is distributed amongst fewer tokens than $K$.
Nucleus Sampling, instead of considering the Top-$K$ most probable tokens in vocabulary $V$, only considers the smallest set of tokens that have a cumulative likelihood greater than some threshold $p_{i}$. Therefore, each decoding step will consider the most likely Top-$p_{i}$ tokens $V^{(p_{i})}\subset V$, automatically eliminating improbable tokens from being accidentally sampled through masking (Holtzman et al. 2020).
@@ A.5 What is a Multi-Agent Simulation?
In a Multi-Agent simulation, independent agents collaborate and interact to find a solution (Russell and Norvig 2009). In the context of our experiments, we have two types of agents: the Query Perturbator and $n+1$ Output Generators. The purpose of the Multi-Agent simulation is to generate independent outputs after perturbing the query $q_{0}$ to observe the rate of hallucination. The presence of $n+1$ queries and agents is to ensure a balanced representation of hallucination, preventing skewed results towards either extreme. The set of outputs generated by the agents is then evaluated against the ground truth value $y$ to estimate the hallucination rate. Through multiple samplings across all 369,837 queries, we train HalluciBot to accurately identify and assess the risk of hallucination.
@@ A.6 What is Monte Carlo Sampling?
Monte Carlo sampling is a technique used to approximate unknown quantities of interest when it’s difficult to derive an exact solution (Murphy 2012). This is particularly useful when dealing with hidden latent variables that cannot be directly observed, such as the case of hallucination caused by interacting with a complex random variable (e.g., a Large Language Model). To address this challenge, we use Monte Carlo sampling. The idea is to conduct a Multi-Agent simulation for each query in our training set. By doing so, we can estimate the probability of hallucination by observing the accuracy rate per query across lexical perturbations. The more simulations we conduct, the more accurate our estimation becomes. In our experiments, we sample from a series of 369,837 Multi-Agent simulations that examine hallucination rates across perturbations for any query $q_{0}$. Therefore, we can estimate the hallucination rate $p_{h}(q_{0})$ for query $q_{0}$, given the ground truth $y$ and set of answers $a_{i}\in\mathcal{A}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum\mathbb{I}\big{[}a_{i}\not=y\big{]}$
@@ A.7 What is Ordinal Regression?
Expanding to several classes $K$, we can divide the probability space into $K-1$ cutpoints $c_{1},\dots,c_{K-1}$ with the property that $c_{1}<c_{2}<\cdots c_{K-1}$, such that the probability for each class is defined as follows (Winship and Mare 1984):
@@
To enforce the property $c_{1}<c_{2}<\cdots c_{K-1}$ while allowing our thresholds to be differentiable, we employ a cumulative sum on a set of unbounded, learnable parameters $\theta_{1},\dots,\theta_{K-1}$, transformed by a Softplus ($\beta=1$) to avoid adding negative values; this ensures our thresholds are always increasing (Zheng et al. 2015). Therefore, each cutpoint $c_{k}$ will be of the from:
@@
Appendix B Additional Experiments & Parameters
@@ B.1 LLM Agent Settings
B.2 Aggregated Monte Carlo Results
The aggregated Monte Carlo results for the MULTI approach demonstrate the relative performance of different question-answering scenarios. On average, the results (Table 3) indicate that Extractive outperforms Multiple Choice, which, in turn, outperforms Abstractive when subjected to perturbations. This trend suggests that the performance of gpt-3.5-turbo is influenced by the presence of additional content. Abstractive tasks show the greatest variation in agent response under perturbations, highlighting the effectiveness of added context in mitigating hallucinations (Figures 2 & 4). Table 17 provides a full breakdown.
[…]
    Extractive: With context, gpt-3.5-turbo performs well on the SQuADv2 dataset. The mode accuracy ($91.0\%$) and agreement ($75.3\%$) of the agents are high. Even under radical perturbations, having unaltered context provides robustness to the agent’s capacity to answer correctly.
[…]
    Multiple Choice: Access to answer choices mitigates hallucinations across perturbations. The ensemble accuracy is slightly higher than the baseline ($+0.5\%$), showcasing that multiple agents can (slightly) improve accuracy rates.
[…]
    Abstractive: When no additional context is provided, gpt-3.5-turbo achieves a mode accuracy of $53.9\%$ under perturbations. Interestingly, there is a significant dispersion of hallucination rates compared to other scenarios (Figure 2). Moreover, there is significant variation in results among datasets. For instance, SQuADv2 shows a $-59.0\%$ decrease in baseline accuracy against its Extractive counterpart. In contrast, SciQ benefits in this setting, leading to a $+9.4\%$ increase in mode accuracy, as the likelihood of generating a match increases.
Listing 1: Prompt function used for query rewriting
def prompt_creator(hallucibot_mode=None, hallucibot_outputs=None):
# Naive Rewrite
if hallucibot_mode is None:
preamble = "You are a helpful expert query correction agent designed to improve user queries when needed in a manner that a novice user can easily understand. Your goal is to take an input user query and only rewrite it if you think that a novice user might find it ambigious and fail a downstream task. If it is ambigious rewrite the query while ensuring that all important information in the query given by the user is present in your rewrite. This means that any user who reads the output rewritten by you will be able to easily understand the question and answer it accurately and could cause hallucinations in a downstream task. Note that some questions might already be cleary understood by a novice user and succeed in the downstream task. These questions do not need to be rewritten."
# HalluciBot Feedback
elif hallucibot_mode == "basic":
preamble = f"You are a helpful expert query correction agent designed to improve user queries when needed in a manner that a novice user can easily understand. Your goal is to take an input user query that has been evaluated using a group of expert critics to have a majority label of ‘{hallucibot_outputs}‘ and produce a better rewritten version if needed.\
A label of ‘hallucinate‘ is caused by ambigious text/information in a query that will cause it to be misunderstood by a novice user causing them to fail in a downstream task.\
A label of ‘not hallucinate‘ does not have this issue and can be easily understood by a novice user who will then succeed in the downstream task. If the label is ‘hallucinate‘, your task is to rewrite the query while ensuring that all important information in the query given by the user is present in the your rewrite. This means that any user who reads the output rewritten by you will be able to easily understand the question and answer it accurately.If the label is ‘not hallucinate‘ your task is to return the user input as is without any modifications."
# With Consensus
elif hallucibot_mode == "rbd":
hb_prediction, hb_consensus_prediction = hallucibot_outputs
if hb_consensus_prediction == "LABEL_0":
hb_consensus_prediction = "All of the expert critics returned the same evaluation."
elif hb_consensus_prediction == "LABEL_1":
hb_consensus_prediction = "A minority of expert critics disagreed with the majority critic consensus about the phrase."
preamble = f"You are a helpful expert query correction agent designed to improve user queries when needed in a manner that a novice user can easily understand. Your goal is to take an input user query that has been evaluated using a group of six expert critics to have a majority label of ‘{hb_prediction}‘ and produce a better rewritten version if needed. ‘{hb_consensus_prediction}‘.\
A label of ‘hallucinate‘ is caused by ambigious text/information in a query that will cause it to be misunderstood by a novice user causing them to fail in a downstream task.\
A label of ‘not hallucinate‘ does not have this issue and can be easily understood by a novice user who will then succeed in the downstream task. If the label is ‘hallucinate‘, your task is to rewrite the query while ensuring that all important information in the query given by the user is present in the your rewrite. This means that any user who reads the output rewritten by you will be able to easily understand the question and answer it accurately.If the label is ‘not hallucinate‘ your task is to return the user input as is without any modifications."
suffix = "Carefully read the entire user query and always return the output as a JSON object with just one key labeled ‘rewritten_query‘ that corresponds to the rewritten question. DO NOT ANSWER THE GIVEN USER QUESTION AS THIS WILL AUTOMATICALLY FAIL THE DOWNSTREAM TASK."
[…]
B.3 Token Usage & Statistics
@@
B.4 HalluciBot Training Setup & Metrics
The parameters to our AWS EC2 training instance are provided in Table 7. Additionally, we report our training metrics, including GPU hours and total floating point operations (FLOPS), in Table 11. Finally, the parameters for our backbones are enumerated in Table 12.
B.5 Multi-class Experiments: Learning the expected value of hallucinations
Multi-class Labels.
Multi-class Results.
HalluciBot achieves a validation accuracy of 47.6%, with a Top 3 accuracy of 73.0% for the RoBERTa-large + Scenario model (Table 10).
[…]
Model Train Val Test Train Val Test Train Val Test Train Val Test
+ Scenario 85.5 72.3 67.4 85.5 76.4 69.8 92.5 80.2 77.3 79.5 73.0 63.7
+ Scenario 79.8 73.0 69.0 79.3 76.8 71.7 88.8 81.5 78.4 71.5 72.6 66.0
+ Consensus 79.3 73.0 68.7 79.1 77.0 71.5 87.2 81.0 77.7 71.4 73.3 66.2
+ Calibration 80.3 73.6 69.5 81.4 78.8 73.6 83.6 78.4 75.6 79.2 79.2 71.7
   + $\tau=0.341$ 80.4 73.6 69.5 81.6 80.2 76.0 74.7 72.9 70.3 90.0 89.0 82.6
RoBERTa-large
+ Scenario 84.9 72.9 68.6 85.0 76.9 71.1 92.1 80.9 78.2 78.8 73.2 65.3
+ Consensus 83.9 73.1 68.7 84.0 77.1 71.7 90.6 80.8 77.2 78.3 73.6 66.9
+ Calibration 84.7 73.5 69.2 85.5 78.5 73.0 88.1 78.9 76.1 83.1 78.2 70.1
   + $\tau=0.326$ 84.8 73.6 69.4 83.5 80.0 75.6 75.0 71.8 70.5 94.2 90.4 81.6
Table 9: Full HalluciBot Binary Evaluation Statistics for BERT (Devlin et al. 2018), RoBERTa-base, and RoBERTa-large (Liu et al. 2019). We report the Accuracy, F1, Precision, and Recall for all data splits. Probability threshold $\tau$ is computed along the closed interval $[0,1]$ in increments of $0.001$ to maximize the validation F1 score for the final model. The best ablation per base model is underlined, while the overall best performing model is in bold. For our experiments, the small RoBERTa-base outperformed the BERT and RoBERTa-large models.
[…]
Model Train Val Test Train Val Test Train Val Test
BERT-base-cased 49.6 32.2 24.7 69.7 49.2 40.7 81.4 62.7 56.4
+ Scenario 54.1 38.7 31.3 72.2 54.8 46.1 82.8 67.6 59.3
+ Ordinal 58.7 45.3 38.6 70.0 54.5 48.3 79.0 64.1 59.1
RoBERTa-base 47.6 34.1 26.6 66.2 50.1 42.6 77.9 62.7 57.3
+ Scenario 52.2 41.5 34.4 69.2 57.0 48.4 79.8 68.6 59.5
+ Ordinal 47.8 39.4 37.1 56.7 48.7 46.6 67.0 60.0 57.8
RoBERTa-large
+ Scenario 61.6 47.6 38.8 77.5 62.6 53.1 85.8 73.0 63.8
+ Ordinal 60.8 48.0 40.7 73.6 59.0 52.2 81.9 67.5 62.3
Table 10: HalluciBot Multi-class Evaluation Statistics. Considering the challenge of approximating a random variable and the potential presence of noise in our empirical estimate, we provide accuracy measurements for Top 1, Top 2, and Top 3 predictions.
@@ B.5 Multi-class Experiments: Learning the expected value of hallucinations
Model Size (Hours) FLOP Steps / Second / Second Loss
@@
Model Size (Hours) FLOP Steps / Second / Second Loss
@@
Table 11: HalluciBot training statistics for our binary and multi-class experiments. Size is the number of learnable parameters in the model. Total FLOP is the total number of floating-point operations conducted in the model. Update Steps is the number of parameter updates the Adam optimizer (Kingma and Ba 2017) performs on the model. We trained each model for 5 epochs in total. Training parameters can be found in Table 12. There is no difference in FLOP between BERT and RoBERTa base models as RoBERTa’s parameter increase is concentrated in a larger vocabulary. The Consensus model is fine-tuned using mixed precision (float16), resulting in half the training time. The loss is twice as much in the Consensus models given the dual loss functions for hallucination and consensus labels.
@@
Table 12: Model configuration and hyper-parameters for our HalluciBots. Configurations changed from defaults highlighted in bold. Note that for training batch size and gradient accumulation steps, we only altered roberta-large’s hyper-parameters to fit onto our GPU. All models were trained using HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba 2017). Note: only the Consensus model is tuned in float16, all other models are trained in full precision.